OpenStudy (anonymous):

What is the domain of 1/ x-1 + 1/x+1??

myininaya (myininaya):

\[\frac{1}\frac{x-1}+\frac{1}\frac{x+1}\]

myininaya (myininaya):

\[\frac{1}{x-1}+\frac{1}{x+1}\]

OpenStudy (anonymous):

Yes. ^^

myininaya (myininaya):

look at the first fraction when is that undefined

OpenStudy (anonymous):

When x is -1?

myininaya (myininaya):

the first fraction is undefined when x=1 the second fraction is undefined when x=-1 the domain is all real numbers but x=1 and x=-1 interval notation: (-inf,-1)U(-1,1)U(1,inf)